OpenStudy (anonymous):

Find the sum 6/x-4+3/x+1

OpenStudy (anonymous):

\[\frac{6}{x-4}+\frac{3}{x+1}\] add just like numbers \[\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bc}\]

OpenStudy (anonymous):

\[\frac{6(x+1)+3(x-4)}{(x-4)(x+1)}\] by direct substitution then you can clean up the numerator with the distributive property and combine like terms

OpenStudy (anonymous):

thanks